Grade 12th passMechanics

A block of mass m is placed on a frictionless wedge of mass M whose angle of inclination is 450 . What should be the minimum coeffecient of friction between the wedge and ground so that the wedge doesnt move.

Dheeraj Sappa
8 Years agoGrade 12th pass

1 Answer

Jobin Joy
8 Years ago
mgcos\theta sin\theta /(M+m)g=\mu
This is the answer
Please Verify